OpenStudy (anonymous):

Choose the equation of the line passing through the point (6, -2) and parallel to y = x + 7. y = x + 8 y = x - 4 y = x + 4 y = x - 8

OpenStudy (anonymous):

Why not just go on a online graphing calculator and graph each of the lines to see which one passes through the point and is parallel?

OpenStudy (user917):

2 parralle equations have the same coefficient. In this case it is 1. ( y= 1*x+7) So the equation of the line will be y=x+b. For x=6, y=-2 so b=y-x=-2-6=-8. Answer: y=x-8.